solve the formula A=1/9si for the variable s
Answer by feliz1965(151)

I assume you meant to type A = (1/9)(s).
We can write (1/9)(s) as s/9. Both expressions mean the same thing. To get s/9, I simply multiplied the numerator 1 by s.
We now have the equation: A = s/9
Multiply both sides by 9.
9A = s
=================================
THE PROVE:
A = (1/9)(9A)
A = A...It checks!
The answer for s is 9A.